UT 20131104 LUCI2 Daytime Commissioning

Observer: DThompson (Tucson), JHeidt (remote)
Telescope Operator: GBechetti
LBTO Support:

Summary

Night was partly bad so flexure data in imaging mode
with the N3.75 camera could be taken. A first set was useless
since we had to stop after two elevations since weather cleared up.
Dome as closed again a while later where a full set
of flexure data in three elevations could be derived.

Preparations

Details

20131104 LUCI2 Flexure Calibrations
Have to do them by hand, not scripted.

READ2 GUI panel would not come up when I first logged in.
LUCI2 GUI panel would not close when the X was clicked.
Obs Execution GUI also would not close when the X was clicked.

To get back to a working system:
Stop Readout Manager
Stop GEIRS
Stop Scheduler
Start GEIRS
Start Scheduler
Start Readout Manager
Start from "Observer" icon on desktop

ssh -X readout2@luci
cleanup
start_luci_new -gui

ps -aux | grap java
killall -9 java
Full LUCI2 SW stop/start from LMC

"Open GUI" from LMC for:
Readout Manager
Instrument Manager

Data are in: /dataB/readout2/DATA/20131103
files: luci2.20131103.NNNN.fits

0001 5.0s "dark" (optic sieve + blink_K filters

Telescope to zenith, AZ=120, ROT=-90
Optic Sieve, mirror, N3.75, clear+K

N3.75 imaging Flexure Calibrations
-90 to +450 and back, 30-degree steps

0002 0038 EL=90
0039 0076 EL=55 (#45, 46 duplicate the ROT=90 measurement)
RH dropped, LBT will open for science...
0076 0 EL=20

ROT EL90 EL55
-90 2 39
-60 3 40
-30 4 41
000 5 42
030 6 43
060 7 44
090 8 45,46
120 9 47
150 10 48
180 11 49
210 12 50
240 13 51
270 14 52
300 15 53
330 16 54
360 17 55
390 18 56
420 19 57
450 20 58
420 21 59
390 22 60
360 23 61
330 24 62
300 25 63
270 26 64
240 27 65
210 28 66
180 29 67
150 30 68
120 31 69
090 32 70
060 33 71
030 34 72
000 35 73
-30 36 74
-60 37 75
-90 38 76

Re-closed for weather. Geno called ~4:00AM

Tried 1x5.0s Dark (Optic Sieve + 2 blind filters)
Submit on READ2 GUI does nothing
Closing LMC also does nothing!
Open LMC on different desktop
Stop Readout Manager
Stop GEIRS Service
Start GEIRS Service
Start Readout Manager
Start READ2 GUI

Things seem to be working now.

0077 1x5.0s Dark (Optic Sieve + 2 blind filters)

Change to clear + K
ROT -> -90
0078 - 0086 = -90 to +150, then I missed 180 before moving to 210!
Take a new image...commit in progress, READ2 GUI hung!
...or does it just take a long time to shut down?
0087 - 0096 = -90 to +180, then I missed 210!

Aha...problem of missed frames is that a read of the detector is
sometimes not saving! This is causing the hang of the READ2 GUI.

ROT or CC fault moving to 450. Geno recovered w/o moving rotator, so I am continuing.

started last set at 6:30

ROT EL20 EL55 EL90
-90 97 134 171
-60 98 135 172
-30 99 136 173
000 100 137 174
030 101 138 175
060 102 139 176
090 103 140 177
120 104 141 178
150 105 142 179
180 106 143 180
210 107 144 181
240 108 145 182
270 109 146 183
300 110 147 184
330 111 148 185
360 112 149 186
390 113 150 187
420 114 151 188
450 115 152 189
420 116 153 190
390 117 154 191
360 118 155 192
330 119 156 193
300 120 157 194
270 121 158 195
240 122 159 196
210 123 160 197
180 124 161 198
150 125 162 199
120 126 163 200
090 127 164 201
060 128 165 202
030 129 166 203
000 130 167 204
-30 131 168 205
-60 132 169 206
-90 133 170 207

-- JochenHeidt - 04 Nov 2013
Topic revision: r3 - 06 Nov 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback